What multiplies to give me -18 and adds to give me 3
Answer by stanbon(75887)

What multiplies to give me -18 and adds to give me 3
----
6 and -3
--------------
Cheers,
Stan H.
-----------------
x + y = 3
xy = -18
------
x(3-x) = -18
3x - x^2 = -18
x^2 - 3x - 18 = 0
-------
(x-6)(x+3) = 0
x = 6 or x = -3
---------------------